# Primitive immersions of constant curvature of surfaces into flag manifolds

### v1

@@ 1. Introduction
E. Calabi in his PhD thesis considered the problem of finding which complex Hermitian manifolds can be holomorphically and isometrically embedded into a complex space form (see [25] and the references therein). His answers to this problem lead to important applications, among which we can find the local classification of complex submanifolds with constant holomorphic sectional curvature of a complex space form. In particular, E. Calabi showed that any such submanifold of the complex projective space $\mathbb{C}P^{n-1}$ is locally unitarily congruent to a piece of a round Veronese embedding. Making use of this classification result, J. Bolton et al. [6] proved that any conformal minimal immersion of constant curvature of the two-sphere $S^{2}$ in the complex projetive space belongs to the Veronese sequence, up to unitary congruence. This was later generalized by Q.-S. Chi and Y. Zheng to the case where the domain is a simply connected surface, not necessarily closed, in [10, Theorem 2].
Recall that, for isometric immersions, minimality is equivalent to harmonicity [12, 30]. The study of harmonic maps from Riemann surfaces into complex Grassmanians through their harmonic sequences was initiated by J. G. Wolfson [31] and further developed by J. C. Wood and his collaborators in a sequence of papers [3, 4, 9, 14]. Due to the fact that there does not exist any nonzero holomorphic differentials on $S^{2}$, there are strong orthogonality conditions on harmonic sequences and diagrams associated to harmonic maps from $S^{2}$ into Grassmannians. More recently, the method of harmonic sequences has been intensively used to classify minimal immersions of constant curvature of $S^{2}$ into different Riemannian symmetric spaces [17, 18, 19, 20, 21, 22, 23, 24, 26, 28].
On the other hand, there exists a well-established theory on twistorial constructions of harmonic maps from Riemann surfaces into symmetric spaces [7, 8]. An important class of twistor lifts is that of primitive maps into $k$-symmetric spaces [7, 16].
A natural problem arising from above observations is that of characterizing the primitive immersions with constant curvature of a Riemann surface into a $k$-symmetric space $G/K$ (equipped with $G$-invariant metrics). In the present paper we address this question in the important case of primitive lifts of pseudoholomorphic maps from $S^{2}$ into complex Grassmannians. Recall that, by definition, any pseudoholomorphic map into a complex Grassmannian $G_{r}({\mathbb{C}}^{n})$ belongs to the harmonic sequence $\psi_{0},\ldots,\psi_{p}$ associated to some holomorphic map $\psi_{0}$ (which we will assume to be linearly full) into a complex Grassmannian $G_{r_{0}}({\mathbb{C}}^{n})$, with $r_{0}\geq r$; the corresponding primitive lift $\Psi=(\psi_{0},\ldots,\psi_{p})$ takes values in some flag manifold of the form
@@
with $r_{0}+\ldots+r_{p}=n$, endowed with its natural structure of $k$-symmetric space, with $k=p+1$. Note that the complex Grassmannian admits only one (up to positive scalar multiplication) $\operatorname{U}(n)$-invariant metric, while the flag manifold (for $p>1$) admits infinite $\operatorname{U}(n)$-invariant metrics. In Theorem 2.1 we will prove that if any such primitive lift $\Psi=(\psi_{0},\ldots,\psi_{p})$ from $S^{2}$ has constant curvature with respect to at least one invariant metric, then it has constant curvature with respect to all invariant metrics; moreover, each $\psi_{j}:S^{2}\to G_{k_{j}}({\mathbb{C}}^{n})$ is an immersion of constant curvature with constant Kähler angle. Since all harmonic maps from $S^{2}$ into ${\mathbb{C}}P^{n-1}$ are pseudoholomorphic [13], we will conclude (see Corollary 2.2) that any full primitive immersion from $S^{2}$ into the full flag manifold which has constant curvature with respect to at least one invariant metric is unitarily equivalent to the primitive lift of a Veronese map. In the final section, we prove a partial generalization of this result to the case where the domain is a general simply connected Riemann surface. The technique introduced by Q.-S. Chi and Y. Zheng [10] will play an important role in this generalization. On the way, we consider the problem of finding the invariant metric on the flag manifold, under a certain normalization condition, that maximizes the induced area of $S^{2}$ by a given primitive immersion.
@@ 1.2. Harmonic sequences
We start by recalling from [6, 9, 22, 31] the definition and properties of harmonic sequences associated to harmonic maps from Riemann surfaces into complex Grassmannians.
We consider on ${\mathbb{C}}^{n}$ the standard Hermitian inner product $v,w\right>=v_{1}\overline{w}_{1}+\ldots+v_{n}\overline{w}_{n}.$ The Grassmannian $G_{k}({\mathbb{C}}^{n})$ of all $k$-dimensional complex subspaces of ${\mathbb{C}}^{n}$ is a Hermitian symmetric $\operatorname{U}(n)$-space, with stabilizers conjugate to $\operatorname{U}(k)\times\operatorname{U}(n-k)$. Given $L\in G_{k}({\mathbb{C}}^{n})$, the complex structure at the corresponding tangent space is given by
@@
for $\xi,\eta\in T_{L}G_{k}({\mathbb{C}}^{n})$. When $k=1$, this is the Fubini-Study metric of constant holomorphic sectional curvature equal to 4 on ${\mathbb{C}}P^{n-1}$.
Given a Riemann surface $M$, we interpret any smooth map $\psi:M\to G_{k}({\mathbb{C}}^{n})$ as a (complex) rank-$k$ vector subbundle, also denoted by $\psi$, of the trivial vector bundle $M\times{\mathbb{C}}^{n}$, with fibre at $z\in M$ given by $\psi(z)$. For the theory of harmonic maps from surfaces into complex Grassmanians, we refer the reader to [9, 27].
Let $\psi:M\to G_{k}({\mathbb{C}}^{n})$ be a harmonic map and $\{\psi_{j}\}_{j\in\mathbb{Z}}$ be its harmonic sequence, where: $\psi_{0}=\psi$; for $j>0$, $\psi_{j}$ is the $j$-th $\partial^{\prime}$-Gauss bundle of $\psi$; for $j<0$, $\psi_{j}$ is the $j$-th $\partial^{\prime\prime}$-Gauss bundle of $\psi$ Recall that, if $\psi$ is holomorphic, then its harmonic sequence must take the form
@@
A harmonic map from a Riemann surface into a complex Grassmannian is called pseudoholomorphic if it belongs to the harmonic sequence of some holomorphic map.
@@
    Using the terminology of loop group theory for harmonic maps from Riemann surfaces into symmetric spaces (first introduced by K. Uhlenbeck [29]), all pseudoholomorphic maps are harmonic maps of finite uniton number. General criteria for finiteness of the uniton number for harmonic maps into complex Grassmannians in terms of the corresponding harmonic sequences and diagrams have been recently developed in [1, 27].
@@
    The nonexistence of nonzero holomorphic differentials on the two-sphere $S^{2}$ ensures that all harmonic maps from $S^{2}$ into ${\mathbb{C}}P^{n-1}$ are pseudoholomorphic [13].
Let $\psi_{0}:M\to G_{k}({\mathbb{C}}^{n})$ be a holomorphic map and $\psi_{0},\psi_{1},\dots\psi_{p}$ be its harmonic sequence. Take a local complex coordinate $z$ on $U\subset M$. According to (1), we have, for each $j\in\{0,\ldots,p\}$,
@@
where $A^{\prime}_{\psi_{j}}$ stands for the $\partial^{\prime}$-second fundamental form of $\psi_{j}$. Set
@@
Each $\psi_{j}$ is a branched conformal immersion and the metric induced from (2) by $\psi_{j}$ on $M$ is locally given by $\psi_{j}^{*}h=(\gamma_{j-1}+\gamma_{j})dzd\bar{z};$ the corresponding curvature is given by
@@
Denoting by $\theta_{j}$ the Kähler angle of $\psi_{j}$, we have
@@
Remark 1.2.
Consider the Plücker embedding $\iota:G_{k}({\mathbb{C}}^{n})\hookrightarrow{\mathbb{C}}P^{{n\choose k}-1}.$ This is a holomorphic isometry. Hence, if $\psi:M\to G_{k}({\mathbb{C}}^{n})$ is a holomorphic map, and $s$ is a local nonvanishing section of $\iota\circ\psi$, then the metric on $M$ induced by $\psi$ is locally given by (see [6, 10, 19, 22])
@@ 1.3. The Veronese sequence
Moreover, $V_{j}^{n}$ is a minimal immersion with the induced metric
@@ 1.4. Primitive harmonic immersions
[7] Let $M$ be a Riemann surface. A smooth map $\Psi:M\to F_{k_{0},\ldots,k_{p}}$, with $\Psi=(\psi_{0},\ldots,\psi_{p})$ is said to be primitive if $d\Psi\big{(}\tfrac{\partial}{\partial z}\big{)}$ is a local section of $\Psi^{*}T^{1}$, for all local complex coordinate $z$ on $M$.
@@
    For $p>1$, if $\Psi:M\to F_{k_{0},\ldots,k_{p}}$ is primitive, then each $\psi_{j}:M\to G_{k_{j}}({\mathbb{C}}^{n})$ is harmonic [7]. (For $p=1$, all maps are primitive.)
It is known [5] that, for $p>1$, any primitive map into $F_{k_{0},\ldots,k_{p}}$ is harmonic with respect to all $\operatorname{U}(n)$-invariant metrics of $F_{k_{0},\ldots,k_{p}}$. Any such metric $g$ has the following form [2]: given $\xi,\eta\in T_{\Psi}F_{k_{0},\ldots,k_{p}}$, and writing $\xi=\sum\xi_{ij}$, $\eta=\sum\eta_{ij}$ according to (7), then
@@
for some positive constants $\lambda_{ij}$ satisfying $\lambda_{ij}=\lambda_{ji}$. Hence, while the Grassmannian admits a unique (up to multiplication by a positive constant) $\operatorname{U}(n)$-invariant metric, the flag $F_{k_{0},\ldots,k_{p}}$ admits infinite nonequivalent $\operatorname{U}(n)$-invariant metrics.
@@
[…]
@@ 2. Primitive harmonic immersions of constant curvature
$\Psi^{*}\tilde{g}=\sum_{j=0}^{n-1}\tilde{\lambda}_{j}\gamma_{j}dzd\bar{z}=\sum_{j=0}^{n-1}\frac{\tilde{\lambda}_{j}\alpha_{j}}{(1+z\bar{z})^{2}}dzd\bar{z},$
and the result follows from (10). ∎
@@
Let $\Psi:S^{2}\rightarrow F_{1,\ldots,1}$ be a full primitive immersion. If there exists at least one $\operatorname{U}(n)$-invariant metric on $F_{1,\ldots,1}$ with respect to which $\Psi$ has constant curvature, then $\Psi$ is unitarily congruent with the $(n-1)$-Veronese primitive map.
@@
Since $\Psi$ is a full primitive map, each $\psi_{j}$, with $j\in\mathbb{Z}_{n}$, is a full harmonic map into ${\mathbb{C}}P^{n-1}$ (see Remark 1.5), and the first Gauss bundle of $\psi_{j}$ is either $\{0\}$ or $\psi_{j+1}$. Since all harmonic maps from $S^{2}$ into ${\mathbb{C}}P^{n-1}$ are pseudoholomorphic, we can reorder (this corresponds to a unitarily congruence on the full flag) if necessary, so that $\psi_{0}:S^{2}\to{\mathbb{C}}P^{n-1}$ is holomorphic and $\psi_{0},\ldots,\psi_{n}$ is the corresponding harmonic sequence. By Theorem 2.1, $\psi_{0}$ is an immersion of constant curvature, hence, by Calabi’s result, $\psi_{0}$ is unitarily congruent with the $(n-1)$-Veronese map. ∎
The following example provides evidence for Theorem 2.1.
@@
The nowhere zero local section $f_{0}(z)=(1,az,z^{2})$ satisfies

### v2

@@ 1. Introduction
E. Calabi in his PhD thesis considered the problem of finding which complex Hermitian manifolds can be holomorphically and isometrically embedded into a complex space form (see [25] and the references therein). His answers to this problem led to important applications, among which we can find the local classification of complex submanifolds with constant holomorphic sectional curvature of a complex space form. In particular, E. Calabi showed that any such submanifold of the complex projective space $\mathbb{C}P^{n-1}$ is locally unitarily congruent to a piece of a round Veronese embedding. Making use of this classification result, J. Bolton et al. [6] proved that any conformal minimal immersion of constant curvature of the two-sphere $S^{2}$ in the complex projetive space belongs to the Veronese sequence, up to unitary congruence. This was later generalized by Q.-S. Chi and Y. Zheng to the case where the domain is a simply connected surface, not necessarily closed, in [10, Theorem 2].
For isometric immersions, minimality is equivalent to harmonicity [12, 30], hence the methods of harmonic map theory can be brought to the study of minimal immersions of constant curvature. We are particularly interested in the following two methods:
  1. (1)
    Harmonic sequences. The study of harmonic maps from Riemann surfaces into complex Grassmanians through their harmonic sequences was initiated by J. G. Wolfson [31] and further developed by J. C. Wood and his collaborators in a sequence of papers [3, 4, 9, 14]. More recently, the method of harmonic sequences has been intensively used to classify minimal immersions of constant curvature of $S^{2}$ into different Riemannian symmetric spaces [17, 18, 19, 20, 21, 22, 23, 24, 26, 28].
  2. (2)
    Twistorial constructions. On the other hand, there exists a well-established theory on twistorial constructions of harmonic maps from Riemann surfaces into symmetric spaces [7, 8]. An important class of twistor lifts is that of primitive maps into $k$-symmetric spaces [7, 16].
A natural problem arising from the above observations is that of characterizing the primitive immersions with constant curvature of a Riemann surface into a $k$-symmetric space $G/K$ (equipped with $G$-invariant metrics). In the present paper we address this question in the important case of primitive lifts of pseudoholomorphic maps [22] from $S^{2}$ into complex Grassmannians. Recall that, by definition, any pseudoholomorphic map into a complex Grassmannian $G_{r}({\mathbb{C}}^{n})$ belongs to the harmonic sequence $\psi_{0},\ldots,\psi_{p}$ associated to some holomorphic map $\psi_{0}$ (which we will assume to be linearly full) into a complex Grassmannian $G_{r_{0}}({\mathbb{C}}^{n})$, with $r_{0}\geq r$; the corresponding primitive lift $\Psi=(\psi_{0},\ldots,\psi_{p})$ takes values in some flag manifold of the form
@@
with $r_{0}+\ldots+r_{p}=n$, endowed with its natural structure of $k$-symmetric space, with $k=p+1$. Note that the complex Grassmannian admits only one (up to positive scalar multiplication) $\operatorname{U}(n)$-invariant metric, while the flag manifold (for $p>1$) admits infinitely many nonequivalent $\operatorname{U}(n)$-invariant metrics. In Theorem 3.1 we will prove that if any such primitive lift $\Psi=(\psi_{0},\ldots,\psi_{p})$ from $S^{2}$ has constant curvature with respect to at least one invariant metric, then it has constant curvature with respect to all invariant metrics; moreover, each $\psi_{j}:S^{2}\to G_{r_{j}}({\mathbb{C}}^{n})$ is an immersion of constant curvature with constant Kähler angle. Since all harmonic maps from $S^{2}$ into ${\mathbb{C}}P^{n-1}$ are pseudoholomorphic [13], we will conclude (see Corollary 3.2) that any full primitive immersion from $S^{2}$ into the full flag manifold which has constant curvature with respect to at least one invariant metric is unitarily equivalent to the primitive lift of a Veronese map. In the final section, we prove a partial generalization of this result to the case where the domain is a general simply connected Riemann surface (not necessarily closed). The technique introduced by Q.-S. Chi and Y. Zheng [10] will play an important role in this generalization. On the way, we consider the problem of finding the invariant metric on the flag manifold, under a certain normalization condition, that maximizes the induced area of $S^{2}$ by a given primitive immersion.
@@ 2.1. Harmonic sequences
We start by recalling from [6, 9, 14, 22, 31] the definition and properties of harmonic sequences associated to harmonic maps from Riemann surfaces into complex Grassmannians.
We consider on ${\mathbb{C}}^{n}$ the standard Hermitian inner product
$\left<v,w\right>=v_{1}\overline{w}_{1}+\ldots+v_{n}\overline{w}_{n},$
for $v=(v_{1},\ldots,v_{n}),w=(w_{1},\ldots,w_{n})\in{\mathbb{C}}^{n}$. The Grassmannian $G_{k}({\mathbb{C}}^{n})$ of all $k$-dimensional complex subspaces of ${\mathbb{C}}^{n}$ is a Hermitian symmetric $\operatorname{U}(n)$-space, with stabilizers conjugate to $\operatorname{U}(k)\times\operatorname{U}(n-k)$. Given $L\in G_{k}({\mathbb{C}}^{n})$, the complex structure at the corresponding tangent space is given by [9, §0.(B)]
@@
for $\xi,\eta\in T_{L}G_{k}({\mathbb{C}}^{n})$. When $k=1$, this is the Fubini-Study metric of constant holomorphic sectional curvature equal to 4 on ${\mathbb{C}}P^{n-1}$ (see also [6, §2.]).
Given a Riemann surface $M$, we interpret any smooth map $\psi:M\to G_{k}({\mathbb{C}}^{n})$ as a (complex) rank-$k$ vector subbundle, also denoted by $\psi$, of the trivial vector bundle $M\times{\mathbb{C}}^{n}$, with fibre at $z\in M$ given by $\psi(z)$.
Let $\psi:M\to G_{k}({\mathbb{C}}^{n})$ be a smooth map. Take a local complex coordinate $z$ on $U\subset M$. The second fundamental forms $A^{\prime}_{\psi},A^{\prime\prime}_{\psi}\in\operatorname{Hom}(\psi|_{U},\psi^{\perp}|_{U})$ (see [9, §1.]) are defined by
$A^{\prime}_{\psi}(s)=\pi_{\psi^{\perp}}\circ\frac{\partial}{\partial z}s,\quad A^{\prime\prime}_{\psi}(s)=\pi_{\psi^{\perp}}\circ\frac{\partial}{\partial\bar{z}}s,$
where $s$ is a section of $\psi|_{U}$ and $\pi_{\psi^{\perp}}$ is the orthogonal (Hermitian) projection onto $\psi^{\perp}$. We clearly have $A^{\prime\prime}_{\psi}=-(A^{\prime}_{\psi^{\perp}})^{*}$. These objects can be turned into global objects by considering the vector bundle homomorphisms $\mathcal{A}^{\prime}_{\psi}\in\operatorname{Hom}(T^{1,0}M\otimes\psi,\psi^{\perp})$ and $\mathcal{A}^{\prime\prime}_{\psi}\in\operatorname{Hom}(T^{0,1}M\otimes\psi,\psi^{\perp})$ given locally by ${A}^{\prime}_{\psi}\otimes dz$ and ${A}^{\prime\prime}_{\psi}\otimes d\bar{z}$, respectively. It is well known [9, Lemma 1.3] that $\psi$ is harmonic if and only if $\mathcal{A}^{\prime}_{\psi}$ is holomorphic, and this holds if and only if $\mathcal{A}^{\prime\prime}_{\psi}$ is antiholomorphic (with respect to the Koszul-Malgrange complex structures of $\psi$ and $\psi^{\perp}$).
Given a harmonic map $\psi:M\to G_{k}({\mathbb{C}}^{n})$, the holomorphicity of $\mathcal{A}^{\prime}_{\psi}$ is used (see [9, §2.]) to define the first $\partial^{\prime}$-Gauss bundle $\psi_{1}=\underline{\operatorname{Im}}\,\mathcal{A}_{\psi}^{\prime}$ of $\psi$ as the unique vector subbundle of $\psi^{\perp}$ that coincides with the image of $\mathcal{A}_{\psi}^{\prime}$ almost everywhere. If $\psi_{1}=\{0\}$, then $\psi$ is antiholomorphic; otherwise $\psi_{1}$ is a harmonic subbundle of rank $k^{\prime}\leq k$. Similarly, the antiholomorphicity of $\mathcal{A}^{\prime\prime}_{\psi}$ is used to define the first $\partial^{\prime\prime}$-Gauss bundle $\psi_{-1}=\underline{\operatorname{Im}}\,\mathcal{A}_{\psi}^{\prime\prime}$ of $\psi$. If $\psi_{-1}=\{0\}$, then $\psi$ is holomorphic; otherwise $\psi_{-1}$ is a harmonic subbundle of rank $k^{\prime\prime}\leq k$. Hence, proceeding recursively, we can associate to a harmonic map $\psi:M\to G_{k}({\mathbb{C}}^{n})$ a sequence $\{\psi_{j}\}_{j\in\mathbb{Z}}$ of harmonic subbundles, with $\psi=\psi_{0}$, which is called the associated harmonic sequence (see [9, §2. and §3.]).
If $\psi$ is holomorphic, then $\psi$ is strongly isotropic [14, (1.7)] (i.e. it has isotropy order $\geq r$ for all $r$); it follows from [9, Lemma 3.1] that the harmonic sequence of $\psi$ must take the form
@@
A harmonic map from a Riemann surface into a complex Grassmannian is called pseudoholomorphic if it belongs to the harmonic sequence of some holomorphic map. In particular, all pseudoholomorphic maps are strongly isotropic.
@@
    Using the terminology of loop group theory for harmonic maps from Riemann surfaces into symmetric spaces (first introduced by K. Uhlenbeck [29]), all pseudoholomorphic maps are harmonic maps of finite uniton number [1, Proposition 4.9]. General criteria for finiteness of the uniton number for harmonic maps into complex Grassmannians in terms of the corresponding harmonic sequences and diagrams have been recently developed in [1, 27].
@@
    The nonexistence of nonzero holomorphic differentials on the two-sphere $S^{2}$ ensures that all harmonic maps from $S^{2}$ into ${\mathbb{C}}P^{n-1}$ are pseudoholomorphic (see [9, Theorem 2.6] and [13, §7.(C)]).
  3. c)
    An example of a harmonic map from the two-torus $T^{2}$ into a complex projective space which is not pseudoholomorphic is provided by the Clifford solution $\psi:T^{2}\to{\mathbb{C}}P^{3}$, which is defined in homogeneous coordinates by
    $\psi(z)=[e^{2i(z+\bar{z})},ie^{-2i(z+\bar{z})},e^{2(\bar{z}-z)},e^{2(z-\bar{z})}].$
Let $\psi_{0}:M\to G_{k}({\mathbb{C}}^{n})$ be a holomorphic map and $\psi_{0},\psi_{1},\dots,\psi_{p}$ be its harmonic sequence. In view of (1), we have (see [9, §1.C])
@@
for each $j\in\{0,\ldots,p\}$. Set
@@
Each $\psi_{j}$ is a branched conformal immersion and the metric induced from (2) by $\psi_{j}$ on $M$ is locally given by $\psi_{j}^{*}h=(\gamma_{j-1}+\gamma_{j})dzd\bar{z};$ the corresponding curvature is given by [22, §3.]
@@
Denoting by $\theta_{j}$ the Kähler angle of $\psi_{j}$, we have [22, §3.]
@@
Remark $2.$
Consider the Plücker embedding $\iota:G_{k}({\mathbb{C}}^{n})\hookrightarrow{\mathbb{C}}P^{{n\choose k}-1}.$ This is a holomorphic isometry. Hence, if $\psi:M\to G_{k}({\mathbb{C}}^{n})$ is a holomorphic map, and $s$ is a local nonvanishing section of $\iota\circ\psi$, then the metric on $M$ induced by $\psi$ from $h=ds^{2}_{G_{k}({\mathbb{C}}^{n})}$ is locally given by (see [6, 10, 19, 22])
@@ 2.2. The Veronese sequence
Moreover, $V_{j}^{n}$ is a minimal immersion with induced metric
@@ 2.3. Primitive harmonic immersions
[7, §1.3] Let $M$ be a Riemann surface. A smooth map $\Psi:M\to F_{k_{0},\ldots,k_{p}}$, with $\Psi=(\psi_{0},\ldots,\psi_{p})$ is said to be primitive if $d\Psi\big(\tfrac{\partial}{\partial z}\big)$ is a local section of $\Psi^{*}T^{1}$, for all local complex coordinate $z$ on $M$.
@@
    For $p>1$, if $\Psi:M\to F_{k_{0},\ldots,k_{p}}$ is primitive, then each $\psi_{j}:M\to G_{k_{j}}({\mathbb{C}}^{n})$ is harmonic [7, Proposition 1.3]. (For $p=1$, all maps are primitive.)
A result due to M. Black [5] (see [7, Proposition 1.1]) implies that, for $p>1$, any primitive map into $F_{k_{0},\ldots,k_{p}}$ is harmonic with respect to all $\operatorname{U}(n)$-invariant metrics of $F_{k_{0},\ldots,k_{p}}$. Any such metric $g$ has the following form [2, p.104]: given $\xi,\eta\in T_{\Psi}F_{k_{0},\ldots,k_{p}}$, and writing $\xi=\sum\xi_{ij}$, $\eta=\sum\eta_{ij}$ according to (7), then
@@
for some positive constants $\lambda_{ij}$ satisfying $\lambda_{ij}=\lambda_{ji}$. Hence, while the Grassmannian admits a unique (up to multiplication by a positive constant) $\operatorname{U}(n)$-invariant metric, the flag $F_{k_{0},\ldots,k_{p}}$ admits infinitely many nonequivalent $\operatorname{U}(n)$-invariant metrics.
@@
We consider the $n$-Veronese sequence $V^{n}_{0},\ldots,V_{n}^{n}$ and the corresponding primitive map $\mathcal{V}=(V^{n}_{0},\ldots,V_{n}^{n}):S^{2}\to F_{1,\ldots,1}$, which will be called the $n$-Veronese primitive map. In view of Theorem 2.3, given an invariant metric on $F_{1,\ldots,1}$ with positive constants $\lambda_{0},\ldots,\lambda_{n-1}$, according to (9), the induced metric on $S^{2}$ from the flag is given by
@@ 3. Primitive harmonic immersions of constant curvature
$\Psi^{*}\tilde{g}=\sum_{j=0}^{n-1}\tilde{\lambda}_{j}\gamma_{j}dzd\bar{z}=\frac{\sum_{j=0}^{n-1}\tilde{\lambda}_{j}\alpha_{j}}{(1+z\bar{z})^{2}}dzd\bar{z}.$
Using (10) to compute the curvature, it follows that $\Psi$ has constant curvature $\frac{4}{\sum_{j=0}^{n-1}\tilde{\lambda}_{j}\alpha_{j}}$ with respect to $\tilde{g}$
@@
Let $\Psi:S^{2}\rightarrow F_{1,\ldots,1}$ be a full primitive immersion. If there exists at least one $\operatorname{U}(n)$-invariant metric on $F_{1,\ldots,1}$ with respect to which $\Psi$ has constant curvature, then $\Psi$ is unitarily congruent with the (n1)(n-1)-Veronese primitive map.
@@
Since $\Psi=(\psi_{0},\ldots,\psi_{n-1})$ is a full primitive map, each $\psi_{j}$, with $j\in\{0,\ldots,n-1\}$, is a full harmonic map into ${\mathbb{C}}P^{n-1}$ (see Remark 2.5), and the first Gauss bundle of $\psi_{j}$ is either $\{0\}$ or $\psi_{j+1}$. Since all harmonic maps from $S^{2}$ into ${\mathbb{C}}P^{n-1}$ are pseudoholomorphic, we can reorder (this corresponds to a unitarily congruence on the full flag) if necessary, so that $\psi_{0}:S^{2}\to{\mathbb{C}}P^{n-1}$ is holomorphic and $\psi_{0},\ldots,\psi_{n}$ is the corresponding harmonic sequence. By Theorem 3.1, $\psi_{0}$ is an immersion of constant curvature, hence, by Calabi’s result, $\psi_{0}$ is unitarily congruent with the (n1)(n-1)-Veronese map. ∎
@@
Observe that $\psi_{0}^{\sqrt{2}}=V^{2}$. The nowhere zero local section $f_{0}(z)=(1,az,z^{2})$ satisfies